# CompMarkGS: Robust Watermarking for Compressed 3D Gaussian Splatting

### v5

@@ 4.3 Frequency-aware anchor growing
We apply a Discrete Wavelet Transform (DWT) to the rendered image and feed its low-frequency Low-Low ($LL$) subband, $I_{LL}$, into the pre-trained HiDDeN zhu2018hidden decoder $D$ to extract the watermark message $M^{\prime}$, which is defined as follows:
$M^{\prime}=D(I_{LL}).$ (5)
However, since the low-frequency band encodes the global structure of the image, embedding a watermark in this region may induce subtle deformations of the overall structure, resulting in blurred boundaries or slight texture distortions. The conventional anchor-growing strategy generates new anchors in voxels where Gaussian gradients are large during training, distributing anchors uniformly across the scene without regard to specific regions, which can cause loss of fine detail.
To address this, we propose a frequency-aware anchor growing strategy that selectively grows anchors in high-frequency regions rich in detail, such as textures. As shown in Fig. 2, we first transform the rendered image $I^{\prime}$ and the ground truth image $I$ into the frequency domain via the Discrete Fourier Transform (DFT). We then apply a high-pass filter to isolate the high-frequency components and reconstruct the high-frequency images $I^{\prime}_{hf}$, and $I_{hf}$ by applying the Inverse Discrete Fourier Transform (IDFT). To generate a binary mask that selects pixels in the high-frequency region, we define the pixel-wise SSIM loss $P_{error}$ as follows:
@@
where $p$ denotes a pixel coordinate in the image, $I^{\prime}_{hf}(p)$ denotes the pixel value in the high-frequency region of the rendered image, and $I_{hf}(p)$ corresponds to a pixel value in the high-frequency region of the ground truth image. After computing the median of all pixel-wise SSIM losses, denoted as $\tilde{P}_{error}$, we select pixels within the range $[\tilde{P}_{error}-\epsilon,\tilde{P}_{error}+\epsilon]$ to create a binary mask, which is then used to extract the 2D coordinates of pixels in the high-frequency region. Subsequently, each 3D Gaussian point is projected onto the 2D image plane, and its continuous coordinates are rounded to obtain integer pixel-level coordinates. We match the 2D Gaussian coordinates with the high-frequency pixel coordinates, generating a boolean mask $F_{mask}$ that identifies anchors located in the high-frequency region. The Gaussians selected by $F_{mask}$ are considered active for the anchor growing process.
@@ 4.4 Objective function
We incorporate the HSV loss $\mathcal{L}_{hsv}$ term to minimize color artifacts during watermark embedding. Previous studies huang2025gaussianmarker ; jang2024waterf ; jang20243dgsw3dgaussiansplatting typically computed image loss in the RGB space to improve rendering quality. However, since the RGB space calculates loss based on independent pixel-level color information, it does not accurately reflect the human visual system. Therefore, we compute image loss in the HSV space, which represents colors in a manner more similar to the human visual system. First, we construct binary masks based on hue ranges for each color to localize color artifacts. Binary mask is defined as follows:
@@
where $H_{c}$ denotes the hue range corresponding to color $c\in\mathcal{C}$, with $\mathcal{C}=\{R,G,B\}$. $R$, $G$, and $B$ represent the red, green, and blue color channels, respectively.
Then, HSV loss $\mathcal{L}_{\text{hsv}}$ is computed as the mean squared error (MSE) between the pixel value of rendered image $I(p)$ and the ground truth $I_{gt}(p)$ over the pixels in the set $\Omega$ where $M_{c}(p)=1$:
@@
The overall color loss is defined as the average of the losses for each target color in the set $\mathcal{C}$.
Watermark message loss.
As described in Sec.4.3, we use the pre-trained HiDDeN zhu2018hidden decoder to produce the watermark message $M^{\prime}$, and the watermark message loss $\mathcal{L}_{msg}$ is defined as a binary cross entropy (BCE) loss:
$\mathcal{L}_{msg}=$ $-\sum^{L}_{i=1}\Bigl{[}M_{i}\mathrm{log}(\sigma(M^{\prime}_{i}))+(1-M_{i})\mathrm{log}(1-\sigma(M^{\prime}_{i}))\Bigr{]},$ (9)
where $M\in\{0,1\}^{L}$ is the ground truth message, and a sigmoid function $\sigma$ constrains $M^{\prime}$ to $[0,1]$
@@
where $\mathcal{L}_{freq}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}P_{error}(p)$, as the mean of all pixel-wise SSIM losses. $\mathcal{L}_{scaffold}$ represents the reconstruction loss defined in Scaffold-GS lu2024scaffold , which consists of one scale regularization and two rendering fidelity terms.
@@ 5.1 Experimental Setting
Following previous works huang2025gaussianmarker ; jang20243dgsw3dgaussiansplatting , we evaluate our method using the Blender mildenhall2021nerf , LLFF mildenhall2019local , and Mip-NeRF 360 barron2022mip datasets. We use all 25 scenes: 8 synthetic bounded scenes from the Blender dataset, 8 forward-facing real-world scenes from the LLFF dataset, and 9 bounded real-world scenes from the Mip-NeRF 360 dataset. We assess performance using 200 test images from the Blender dataset. For the LLFF and Mip-NeRF 360 datasets, we adopt the same data split strategy as Mip-NeRF 360.
@@
Our method trains end-to-end on a single A6000 GPU. We conduct experiments for watermark bit-lengths of 32, 48, and 64, focusing on the 48-bit case, where we evaluate both before and after compression performance. For the decoder, we use a pre-trained HiDDeN zhu2018hidden decoder for each bit-length and keep its parameters fixed during watermark training. We use the following parameters: $\lambda_{img}=10,\lambda_{hsv}=0.6,\lambda_{freq}=0.1,\lambda_{msg}=0.45$. For the quantization scale and anchor growing range in high-frequency regions, we set $\mathrm{Q}_{0}=1,\epsilon=0.3$.
@@
Methods Bit Accuracy(%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ Size(MB) ↓ HAC chen2024hac + WateRF jang2024waterf 91.02 / 54.40 27.36 / 13.63 0.850 / 0.457 0.174 / 0.574 207.72 / 13.66 HAC chen2024hac + GaussianMarker huang2025gaussianmarker 92.00 / 58.34 27.05 / 13.54 0.840 / 0.460 0.193 / 0.571 341.30 / 24.33 HAC chen2024hac + 3D-GSW jang20243dgsw3dgaussiansplatting 90.96 / 53.48 19.57 / 13.13 0.628 / 0.439 0.295 / 0.572 173.64 / 12.93 HAC chen2024hac + CompMarkGS 95.95 / 95.92 27.68 / 27.65 0.856 / 0.852 0.171 / 0.177 208.96 / 12.23 ContextGS wang2024contextgs + WateRF jang2024waterf 92.01 / 90.36 26.64 / 26.47 0.843 / 0.832 0.183 / 0.185 219.48 /   9.88 ContextGS wang2024contextgs + GaussianMarker huang2025gaussianmarker 91.24 / 87.95 26.89 / 26.54 0.839 / 0.827 0.195 / 0.200 342.38 / 17.86 ContextGS wang2024contextgs + 3D-GSW jang20243dgsw3dgaussiansplatting 88.28 / 79.75 19.50 / 19.83 0.627 / 0.617 0.294 / 0.299 177.19 /   9.24 ContextGS wang2024contextgs + CompMarkGS 94.36 / 94.03 27.60 / 27.55 0.845 / 0.844 0.172 / 0.173 73.39 /   5.72
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$ = 0.1) Rotation (±$\pm\pi/6$) Scaling (75%) Gaussian Blur ($\sigma$ = 0.1) Crop (40%) JPEG Compression (50% quality) HAC chen2024hac + WateRF jang2024waterf 91.02 / 54.40 89.75 / 55.90 84.92 / 56.16 86.17 / 55.38 89.33 / 56.34 87.00 / 55.30 84.00 / 57.47 HAC chen2024hac + GaussianMarker huang2025gaussianmarker 92.00 / 58.34 59.08 / 58.33 58.75 / 58.33 58.33 / 58.33 59.08 / 58.33 61.42 / 58.17 58.25 / 58.33 HAC chen2024hac + 3D-GSW jang20243dgsw3dgaussiansplatting 90.96 / 53.48 89.50 / 52.50 86.67 / 52.25 86.42 / 52.08 86.67 / 52.25 86.92 / 52.50 83.83 / 53.25 HAC chen2024hac + CompMarkGS 95.95 / 95.92 96.00 / 95.50 92.25 / 92.08 91.83 / 91.67 95.83 / 95.58 90.67 / 90.83 87.75 / 88.83 ContextGS wang2024contextgs + WateRF jang2024waterf 92.01 / 90.36 92.25 / 91.50 88.75 / 88.00 89.58 / 88.83 92.75 / 91.50 89.50 / 88.25 86.41 / 84.58 ContextGS wang2024contextgs + GaussianMarker huang2025gaussianmarker 91.24 / 87.95 59.08 / 58.92 59.00 / 58.75 58.25 / 58.17 59.17 / 58.92 61.08 / 60.67 58.25 / 58.25 ContextGS wang2024contextgs + 3D-GSW jang20243dgsw3dgaussiansplatting 88.28 / 79.25 87.92 / 80.67 85.08 / 78.75 84.42 / 77.08 88.00 / 81.42 84.33 / 78.50 81.58 / 72.58 ContextGS wang2024contextgs + CompMarkGS 94.36 / 94.03 93.04 / 92.98 90.17 / 90.33 90.46 / 90.06 92.96 / 92.81 89.63 / 89.29 85.21 / 85.31
@@
We perform message extraction experiments before and after compression using anchor-based 3DGS with quantization-based compression, validating the effectiveness of our method. For comparison, we evaluate against the following three baselines: 1) WateRF jang2024waterf : Embedding watermark into NeRF mildenhall2021nerf via frequency-domain (DWT) techniques; 2) GaussianMarker huang2025gaussianmarker : Integrating watermark-specific Gaussians into the 3D Gaussian Splatting (3DGS) kerbl20233d based on per-Gaussian uncertainty estimates; 3) 3D-GSW jang20243dgsw3dgaussiansplatting : Pruning only those Gaussians whose removal minimally impacts rendering quality, embedding the watermark into the remaining primitives. All three baselines originally designed for NeRF or 3DGS were adapted to the anchor-based 3DGS framework for our experiments. We compare their performance before and after compression using two quantization-based compression schemes: HAC chen2024hac and ContextGS wang2024contextgs .
@@
We evaluate the watermark performance of CompMarkGS and the baselines according to three aspects: 1) Fidelity: We measure PSNR, SSIM, and LPIPS zhang2018unreasonable to assess the rendering quality of watermarked models before and after compression by comparing rendered images to the original images. 2) Robustness: We evaluate bit accuracy under various distortions before and after compression. We evaluate performance under distortions for rendered images, including Gaussian noise, rotation, scaling, Gaussian blur, cropping, and JPEG compression. We assess robustness against Gaussian noise, cloning, and pruning attacks for watermarked models. 3) Capacity: We measure bit accuracy for 32, 48, and 64-bit message lengths.
@@ 5.2 Experimental results
We compare rendering quality and bit accuracy before and after compression against baseline methods. The compression technique compresses anchor-based 3DGS models using quantization and entropy encoding. As shown in Tab.1, our proposed method exhibits strong robustness to compression. While some baselines achieve higher rendering quality before compression, conventional methods without the quantization distortion layer suffer significant watermark and scene information degradation after compression (see Fig.3). In particular, the HAC chen2024hac -based method shows the most significant performance drop, as existing approaches fine-tune without accounting for both the quantization process and the interpolation required for HAC entropy encoding.
Table 3: Robustness to model distortion. Based on 48-bit, results are averaged over three datasets with both HAC&\&&ContextGS before (left) and after (right) compression.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$=0.005) Clone (50%) Prune (20%) WateRF jang2024waterf 91.52 / 72.38 81.70 / 63.27 90.23 / 67.45 87.81 / 66.25 GaussianMarker huang2025gaussianmarker 91.62 / 73.15 73.01 / 65.07 89.83 / 72.17 87.43 / 85.73 3D-GSW jang20243dgsw3dgaussiansplatting 89.62 / 62.89 78.56 / 57.72 88.26 / 62.14 84.84 / 59.76 CompMarkGS 95.16 / 94.98 84.58 / 80.10 94.35 / 94.16 93.53 / 93.42
Table 4: Comparison of bit accuracy for our method and baselines at 32, 48, and 64-bit. Results are averaged over three datasets with both HAC&\&&ContextGS before compression.
Bit Accuracy(%) $\uparrow$ Methods 32 bits 48 bits 64 bits WateRF jang2024waterf 92.46 91.52 88.21 GaussianMarker huang2025gaussianmarker 95.07 91.62 79.81 3D-GSW jang20243dgsw3dgaussiansplatting 93.00 89.62 86.31 CompMarkGS 96.52 95.16 91.29
Similarly, the ContextGS wang2024contextgs -based method experiences a decline in bit accuracy and rendering quality. Moreover, 3D-GSW jang20243dgsw3dgaussiansplatting , which performs pruning based on rendering contributions before watermark training, removes both anchors and their associated Gaussians, significantly degrading rendering quality. These results confirm that our proposed watermarking technique is more suited to model compression than existing methods.
@@
The number of embedded watermark bits in the rendered image represents a trade-off between image quality and bit accuracy. We conduct experiments using 32, 48, and 64-bit. As shown in Tab.4, increasing the number of embedded watermark bits results in a degradation of image quality and bit accuracy. However, our proposed method shows that even with a larger number of bits, the performance degradation is reduced compared to others.

### v6

@@ 4.3 Frequency-aware anchor growing
We apply a Discrete Wavelet Transform (DWT) to the rendered image and feed its low-frequency Low-Low ($LL$) subband, $I_{LL}$, into the pre-trained HiDDeN (Zhu et al., 2018) decoder $D$ to extract the watermark message $M^{\prime}=D(I_{LL})$. The low-frequency band is ideal for embedding watermarks to ensure robustness, but it is also critical for overall quality as it contains the image’s global structure. This creates a fundamental conflict. If we attempt to embed a watermark and enhance quality in the same low-frequency band, the quality enhancement algorithm may treat the watermark as unwanted noise and remove it, thereby degrading the watermark’s robustness.
Based on this insight, we propose a frequency-aware anchor growing (FAG) strategy. To selectively grow anchors in detail-rich, high-frequency regions, our method first transforms the rendered image $I^{\prime}$ and the ground truth image $I$ into the frequency domain using the Discrete Fourier Transform (DFT), as shown in Fig. 2. We then apply a high-pass filter to isolate the high-frequency components and reconstruct the corresponding high-frequency images $I^{\prime}_{hf}$, and $I_{hf}$ using the Inverse Discrete Fourier Transform (IDFT). To generate a binary mask that selects pixels in these high-frequency regions, we define the pixel-wise SSIM loss $P_{error}$ as follows:
@@
where $p$ is a pixel coordinate, $I^{\prime}_{hf}(p)$ is the pixel value in the high-frequency region of the rendered image, and $I_{hf}(p)$ is the corresponding value in the ground truth. After computing the median of all pixel-wise SSIM losses, denoted as $\tilde{P}_{error}$, we create a binary mask by selecting pixels within the range $\tilde{P}_{error}-\epsilon,\tilde{P}_{error}+\epsilon]$. This mask is then used to extract the 2D coordinates of high-frequency regions. Subsequently, each 3D Gaussian is projected onto the 2D image plane, and its continuous coordinates are rounded to obtain integer pixel‐level coordinates. By matching these 2D Gaussian coordinates with the high-frequency pixel coordinates, we generate a boolean mask $F_{mask}$ that identifies anchors located in the high-frequency regions. Only the Gaussians selected by $F_{mask}$ are activated for the anchor growing process. More implementation details are in Appendix C.4C.5.
@@ 4.4 Objective function
Watermark message loss.
As described in Sec.4.3, we use the pre-trained HiDDeN (Zhu et al., 2018) decoder to produce the watermark message $M^{\prime}$, and the watermark message loss $\mathcal{L}_{msg}$ is defined as a binary cross entropy (BCE) loss:
$\mathcal{L}_{msg}=\$$-\sum^{L}_{i=1}\Bigl[M_{i}\mathrm{log}(\sigma(M^{\prime}_{i}))+(1-M_{i})\mathrm{log}(1-\sigma(M^{\prime}_{i}))\Bigr],$ (6)
where $M\in\{0,1\}^{L}$ is the ground truth message, and a sigmoid function $\sigma$ constrains $M^{\prime}$ to $[0,1]$.
@@
We incorporate the HSV loss $\mathcal{L}_{hsv}$ term to minimize color artifacts during watermark embedding. Previous studies (Huang et al., 2025; Jang et al., 2024; 2025) typically computed image loss in the RGB space to improve rendering quality. However, since the RGB space calculates loss based on independent pixel-level color information, it does not accurately reflect the human visual system. Therefore, we compute image loss in the HSV space, which represents colors in a manner more similar to the human visual system. First, we construct binary masks based on hue ranges for each color to localize color artifacts. Binary mask is defined as follows:
@@
where $H_{c}$ denotes the hue range corresponding to color $c\in\mathcal{C}$, with $\mathcal{C}=\{R,G,B\}$. $R$, $G$, and $B$ represent the red, green, and blue color channels, respectively. Then, HSV loss $\mathcal{L}_{\text{hsv}}$ is computed as the mean squared error (MSE) between the pixel value of rendered image $I(p)$ and the ground truth $I_{gt}(p)$ over the pixels in the set $\Omega$ where $M_{c}(p)=1$:
@@
The overall color loss is defined as the average of the losses for each target color in the set $\mathcal{C}$. More implementation details are in Appendix C.6.
@@
where $\mathcal{L}_{freq}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}P_{error}(p)$, as the mean of all pixel-wise SSIM losses. $\mathcal{L}_{scaffold}$ represents the reconstruction loss defined in Scaffold-GS (Lu et al., 2024), which consists of one scale regularization and two rendering fidelity terms.
@@ 5.1 Experimental Setting
Following previous works Huang et al. (2025); Jang et al. (2025), we evaluate our method using the Blender Mildenhall et al. (2021), LLFF Mildenhall et al. (2019), and Mip-NeRF 360 Barron et al. (2022) datasets, comprising a total of 25 scenes. These include 8 synthetic bounded scenes from Blender, 8 forward-facing real-world scenes from LLFF, and 9 bounded real-world scenes from Mip-NeRF 360. We evaluate performance using 200 test images from Blender. For the LLFF and Mip-NeRF 360 datasets, we adopt the same data split strategy as Mip-NeRF 360.
@@
Our method trains end-to-end on a single A6000 GPU. We conduct experiments for watermark bit-lengths of 32, 48, and 64, focusing on the 48-bit case, where we evaluate performance both before and after compression. For the decoder, we use a pre-trained HiDDeN (Zhu et al., 2018) decoder for each bit-length and keep its parameters fixed during watermark training. We use the following parameters: $\lambda_{img}=10,\lambda_{hsv}=0.6,\lambda_{freq}=0.1,\lambda_{msg}=0.45$. For the quantization scale and anchor growing range in high-frequency regions, we set $\mathrm{Q}_{0}=1,\epsilon=0.3$. More implementation details are in Appendix C.3.
@@
Methods Bit Accuracy(%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ Size(MB) ↓ HAC + WateRF 91.02 / 54.40 27.36 / 13.63 0.850 / 0.457 0.174 / 0.574 207.72 / 13.66 HAC + GaussianMarker 92.00 / 58.34 27.05 / 13.54 0.840 / 0.460 0.193 / 0.571 341.30 / 24.33 HAC + 3D-GSW 90.96 / 53.48 19.57 / 13.13 0.628 / 0.439 0.295 / 0.572 173.64 / 12.93 HAC + GuardSplat 79.77 / 52.72 16.29 / 12.28 0.584 / 0.425 0.417 / 0.609 215.69 / 13.15 HAC + CompMarkGS 95.95 / 95.92 27.68 / 27.65 0.856 / 0.852 0.171 / 0.177 208.96 / 12.23 ContextGS + WateRF 92.01 / 90.36 26.64 / 26.47 0.843 / 0.832 0.183 / 0.185 219.48 /   9.88 ContextGS + GaussianMarker 91.24 / 87.95 26.89 / 26.54 0.839 / 0.827 0.195 / 0.200 342.38 / 17.86 ContextGS + 3D-GSW 88.28 / 79.75 19.50 / 19.83 0.627 / 0.617 0.294 / 0.299 177.19 /   9.24 ContextGS + GuardSplat 73.20 / 67.16 16.90 / 16.86 0.653 / 0.628 0.353 / 0.360 219.82 /   9.33 ContextGS + CompMarkGS 94.36 / 94.03 27.60 / 27.55 0.845 / 0.844 0.172 / 0.173 73.39 /   5.72
Figure 4: Robustness to image distortion. Based on 48-bit, the results are averaged over three datasets.
Table 2: Robustness to model distortion. Based on 48-bit, results are averaged over three datasets with both HAC&ContextGS before (left) and after (right) compression.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$=0.005) Clone (50%) Prune (20%) WateRF 91.52 / 72.38 81.70 / 63.27 90.23 / 67.45 87.81 / 66.25 GaussianMarker 91.62 / 73.15 73.01 / 65.07 89.83 / 72.17 87.43 / 85.73 3D-GSW 89.62 / 62.89 78.56 / 57.72 88.26 / 62.14 84.84 / 59.76 GuardSplat 76.49 / 59.94 65.28 / 55.53 73.76 / 63.20 69.19 / 65.19 CompMarkGS 95.16 / 94.98 84.58 / 80.10 94.35 / 94.16 93.53 / 93.42
@@
To validate our method, we evaluate watermark extraction before and after quantization-based compression in an anchor-based 3DGS framework. For comparison, we evaluate against the following four baselines: 1) WateRF (Jang et al., 2024): Embedding watermark into NeRF (Mildenhall et al., 2021) via frequency-domain (DWT) techniques; 2) GaussianMarker (Huang et al., 2025): Integrating watermark‐specific Gaussians into the 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) based on per-Gaussian uncertainty estimates; 3) 3D-GSW (Jang et al., 2025): Embedding the watermark message by selectively removing Gaussians with low impact on rendering quality; 4) GuardSplat (Chen et al., 2024b): Leveraging Spherical Harmonics (SH) coefficients to embed the message. All baselines originally designed for NeRF or 3DGS are adapted to the anchor-based 3DGS framework for our experiments. Notably, GuardSplat requires specific modifications due to architectural incompatibilities, which are detailed in Appendix D.1. We compare performance before and after compression using two quantization-based compression schemes: HAC (Chen et al., 2024a) and ContextGS (Wang et al., 2024a).
@@
We evaluate the watermark performance of our proposed method, CompMarkGS, and the baselines according to three aspects: 1) Fidelity: We measure PSNR, SSIM, and LPIPS (Zhang et al., 2018) to assess the rendering quality of watermarked models before and after compression by comparing rendered images to the original images. 2) Robustness: We evaluate bit accuracy under various distortions before and after compression. We evaluate performance under distortions for rendered images, including Gaussian noise, rotation, scaling, Gaussian blur, cropping, and JPEG compression. We assess robustness against Gaussian noise, cloning, and pruning attacks for watermarked models. 3) Capacity: We measure bit accuracy for 32, 48, and 64-bit message lengths.
@@ 5.2 Experimental results
[…]
@@
Table 3: Comparison of bit accuracy for our method and baselines at 32, 48, and 64-bit. Results are averaged over three datasets with both HAC&\&ContextGS before compression.
Bit Accuracy(%) $\uparrow$ Methods 32 bits 48 bits 64 bits WateRF 92.46 91.52 88.21 GaussianMarker 95.07 91.62 79.81 3D-GSW 93.00 89.62 86.31 GuardSplat 79.64 76.49 72.83 CompMarkGS 96.52 95.16 91.29
Increasing a watermark message capacity leaves less embedding budget per bit and makes it more vulnerable to noise and compression, which lowers the bit accuracy. We test bit accuracy for message lengths of 32, 48, and 64 bits. As shown in Tab. 3, while the accuracy for all methods decreases with longer messages, our method maintains significantly higher bit accuracy. In stark contrast to other methods, its minimal performance drop even at a 64-bit messages demonstrates our method’s higher effective messages capacity and greater robustness.
Table 4: Ablations of watermark embedding into different anchor attributes. Results are reported for 48-bit messages and averaged over three datasets using HAC and ContextGS after compression.
HAC ContextGS Target Parameter Bit Acc (%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Bit Acc (%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Position 67.91 12.05 0.625 0.350 82.37 18.49 0.632 0.432 Scaling 65.06 19.35 0.615 0.323 64.68 13.40 0.466 0.474 Offsets 94.88 24.61 0.792 0.226 93.98 25.32 0.755 0.322 Anchor Feature (Ours) 95.92 27.65 0.852 0.177 94.03 27.55 0.844 0.173
HAC ContextGS FAG QDL $\mathcal{L}_{hsv}$ Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ 90.67 26.44 0.827 0.196 87.63 25.56 0.812 0.223 $\checkmark$ $\checkmark$ 93.95 27.54 0.849 0.178 90.35 27.15 0.841 0.198 $\checkmark$ $\checkmark$ 90.75 26.75 0.844 0.182 88.09 25.04 0.801 0.229 $\checkmark$ $\checkmark$ 92.57 27.49 0.852 0.179 90.23 27.19 0.841 0.200 $\checkmark$ $\checkmark$ $\checkmark$ 95.92 27.65 0.852 0.177 94.03 27.55 0.844 0.173